OpenStudy (anonymous):

The equation \(k-3x+x^{3} =0\). Find possible values of \(k\) so that the equation has at most \(2\) roots.

OpenStudy (anonymous):

Well, how am I able to solve this since it's a cubic; I haven't done these in a while. Anyone able to give me a hint?

OpenStudy (anonymous):

Sure! Can we assume k is any constant # though?

OpenStudy (anonymous):

Through trial and error?

Parth (parthkohli):

All solutions are constants, aren't they?

OpenStudy (anonymous):

Because if it's a variable this doesn't clean up neatly.

Parth (parthkohli):

Trial-and-error is my preferable choice in the questions I can't solve :)

OpenStudy (anonymous):

Well, I don't think trial and error is allowed. However, I do not know since I haven't done these in a very long time.

OpenStudy (anonymous):

i think the grade of the x contained by k must be even

Parth (parthkohli):

I am learning such questions, but they are quadratics.

OpenStudy (anonymous):

If it was quadratics then discriminant would be used I think.

Parth (parthkohli):

Yep

OpenStudy (anonymous):

I am very rusty in these :(

OpenStudy (anonymous):

no problem a polynome function has most 2 roots, when the highest polynom is even (e.g. 2,4,6,...)

OpenStudy (anonymous):

which means you assume a solution..?

OpenStudy (anonymous):

0 would be possible too or e.g. \[x^2 \]

OpenStudy (anonymous):

but you do not know what is \(x\)

OpenStudy (anonymous):

Explicit way to your answer http://en.wikipedia.org/wiki/Cubic_function Implicit way we must think.....

OpenStudy (anonymous):

sorry k= [x^2\] or k=[x^4\]

OpenStudy (anonymous):

that is a rule

OpenStudy (anonymous):

which means we guess an answer?

OpenStudy (lgbasallote):

variable separable....

OpenStudy (anonymous):

with an uneven polynome function you would get at most so much roots like the number of the highest polynom is

OpenStudy (anonymous):

which basically means assumptions.. I don't you're allowed to assume in these question.. @lgbasallote: any ideas?

OpenStudy (lgbasallote):

nawp. you guys are too smart for me

OpenStudy (anonymous):

lol, this is elementary (:

OpenStudy (anonymous):

\(k+x^3-3 x = 0\) Subtract k from both sides: \(-k = -3x+x^3\) Multiply by (-1) to both sides: \(k = 3x-x^3\) Factor: \(k = -x (x^2-3)\) So \(x = -k\) or \(x^2-3=k\) \(x^2=k+3\) \(x=\pm \sqrt{k+3}\) So... That's 3 roots. But we can solve the roots implicitly by... http://en.wikipedia.org/wiki/Cubic_function#General_formula_of_roots , if you let it be in the form of ax^3+bx^2+cx+d=0, a\(\neq\)0

OpenStudy (anonymous):

K=x^g for g is only an even figure and equal zero. Or am I absolutely wrong. you are starting to confuse

OpenStudy (anonymous):

well, \(k\) is in terms of \(x\) then which is the roots?

OpenStudy (anonymous):

only answers k=2 or -2

OpenStudy (anonymous):

@Omniscience have the answer?

OpenStudy (anonymous):

Nope. :( I will get the answer tomorrow though.

OpenStudy (anonymous):

Arg I can't type fast enough... If k = 2